#                                         SOCCOM Additional Information - 24-May-2018 05:04:58
7567,5904182, 03 Feb 2016: change notation to MPD. Battery has died while float was under ice, was weakening before it went under earlier in 2015

9092,5904185, 22 Jan 2017:  Additional CTD calibration cast from P18,  EXPOCODE=33RO20161119, Station=195, Cast=1

9031,5904396, 17 Jan 2017: appears to have died at half of battery capacity - probably blew 2 fuses

9018,5904186, Apr 2016:  Problem with conductivity, which is also used to calculate oxygen

9095,5904188, 7 Jun 2016: grounded on 800m deep ridge south of Heezen Fracture Zone after 113 profiles
               06 Jul 2016: change notation to MPD.
               29 Nov 2016: change notation to active: no longer stuck in mud, profile 130
               06 Jul 2016: change notation to MPD.
              06 Mar 2018: Batteries Dead 9254,5904395, 2017 Oct 3 - will be recovered, battery prematurely low. Probably recovered about Oct 17 9254,5904395, 2017 Oct 31 - was recovered by this date 0037,5904475, 18 Jan 2016: float stuck at surface starting 26 Dec 2015 profile 69 - air/oil separator valve stuck
              19 July 2016: began profiling again on 9 July 2016 profile 89
              03 Aug 2016: float is responding again, removed the stuck surface designation 9313,5904474, 2 Aug 2016: Now too far into subtropics for standard eMLR algorithm to work properly

13 Jun 2017: pH bad starting cycle 108 0508,5904476, 17 Jan 2015: pH sensor failed on deployment.
              07 Feb 2016: induced into under-ice mode for testing, following profile 42 on January 23, 2016, until end of February, 2016.
              19 Apr 2016: Declared MPD after missing 42 profiles. Problem with maintaining buoyancy, similar failure mode as several other Seabird Navis floats (0509, 0511) from the same deployment group.
              12 Jul 2016: Potential that float grounded at Marion Island, but most likely buoyancy failure as noted in April. 0509,5904477, 2 June 2015: last good profile before ice cover, did not emerge, declared MPD fall 2015 0511,5904478, 17 Jan 2015: Stuck at surface after 3 good profiles, buoyancy problem
              13 June 2015: UW data acquisition terminated 9094,5904471, 07 Feb 2016: pH flag: removed comments from float status page, put comments in additional information file 9275,5904472, 24 Jan 2018: pH bad after cycle 108 9260,5904473, 17 Jan 2017: sending partial files, but not completing the transmissions. Pneumatic system has failed, bladder fails to inflate, so antenna remains submerged too long to communicate. 9260,5904473, 27 Feb 2017 Lost Buoyancy. Can not stay at surface long enough to transmit. 8514,5904470, 2 Aug 2016: pH sensor becoming bad at profile 54 8514,5904470, 10 Oct 2017: Dead after 96 profiles, last profile on 2017-09-10.This was a float that was used in the initial pH sensor tests off Hawaii and was refurbished for deployment in SOCCOM. 9666,5904662, 18 Jan 2016: salinity problems, not likely salvageable (Swift)
              12 July 2016: bad salinity and other properties below 1000 on profiles 2-7, 11-13, 14, 18; salinity spiking on many of these and also 10, 13, 15 9652,5904660, 07 Feb 2016: keep watch on oxygen; sensor cleaned prior to deployment 9652,5904660, 22 Nov 2017: Additional CTD calibration cast from UK COMICS program (and ORCHESTRA), about 14 hours behind the float surface time, and near the surface location. 9657,5904659, 07 Feb 2016: keep watch on oxygen; sensor cleaned prior to deployment 9645,5904676, 7 June 2016: 9 bad nitrate profiles, can deal with in calibration 0506,5904670, 07 Feb 2016: pH flag: removed comments from float status page, put comments in additional information file 0506,5904670, 13 Jan 2017: Dead CTD 0506,5904670, 10 Aug 2017: The root-cause of the salinity failure on navis 0506 is that the float experienced flooding after profile 035. This diagnosis is pretty conclusive as it involves many interlocking bits of evidence & data. The symptoms are also many ranging from buoyancy issues, faulty valve operations, salinity, malfunction of the pneumatic system, poor GPS and telemetry, malfunction of the PPI that services the suna, and on and on. 9600,5904688, 03 May 2016: 6 profiles completed. pH sensor is beginning to become intermittent, in same failure mode as other pH sensors
              28 June 2016: pH sensor continuing with problems 9631,5904677, 27 Apr 2017: Profiles starting at 33, on Jan. 14, 2017, have been in water around 1000 m on the Campbell Plateau. 0568,5904767, 5 May 2016: pH first profiles shakey 0565,5904672, 09 May 2016: pH sensor not working from start, likely due to leak between back of ISFET and seawater
              09 May 2016: Optical sensor has chlorophyll fluorescence and backscatter at 530 and 700 nm instead of chlorophyll and CDOM fluorosence plus BB at 700 nm 9761,5904764, 12 Jul 2016: CApex (carbon fiber hull Apex) float over-ballasted and probably descended below its crush depth on the 2nd profile and imploded 0571,5904673, 17 Jan 2017: pH sensor failed from outset, data are bad 0566,5904766, 16 May 2016: First pH profile with insufficiently conditioned sensor (SBE-built pH sensor), should improve 9660,5904761, 17 May 2016: bottom of float hit side of ship on deployment, no external sensors were involved.
              03 Apr 2018: pH data set is being recalibrated. pH sensor is not shielded from light and the data are being adjusted (affects top 50 m).
              04 Apr 2018: The pH sensor on float 9660 (WMO 5904761) is not within the pumped flow path and is exposed to ambient light which affects the sensor response. This results in artificially high pH values for surface samples (<50m) in daytime profiles (by a magnitude of 0.01 pH units on average for affected samples, and up to 0.1 for some samples). This data has now been corrected in the *QC.TXT files using diagnostic data returned from the pH sensor. Be aware that files downloaded before today (4/4/2018) will not have this correction. A special note now exists within the header of each *QC.TXT file stating the issue and associated correction method. Raw files have not been modified.
               pH_new = pH_old + (Ib - Ibo) * 1.488e6.
               Ib = base current, Ibo = median base current between 70-120 meters 9762,5904765, 12 Jul 2016: Profiles cleaned up and online. Profile 2 has salinity and FLBB backscatter QC problem, possibly something attached to float and then fell off 9659,5904843, 9659,5904843, 12 Apr 2017: O2 sensor bad; pH still looks good; NO3 flagged good even though O2 (used in qc process) is bad 9642,5904685, 27 Dec 2017: pH failed at cycle 37 12382,5904841, 12 Jan 2017: Nitrate sensor light throughput is ~10% of calibration. Cy2 will help determine cause.

Failed NO3 and pH, O2 gain assessed 9750,5904845, 23 Jan 2017: Gouges with paint scrapes on hull prior to deployment, repaired with light enamel and deployed 8501,5905075, 23 Jan 2017: Gouges with paint scrapes on hull prior to deployment, repaired with light enamel and deployed 12559,5904855, 29 Jan 2017: Additional CTD calibration cast from P18, EXPOCODE=33RO20161119, Station=212, Cast=1 12549,5904859, 13 Oct 2017: pH failed at Cy 29 12390,5904981, 03 Apr 2018: CTD failed on profile 42, GPS fix was received and oxygen air calibration on last profile was receved. Likely to have come up in ice during the mandatory February reporting period 12551,5904858, 10 Jan 2016: optode failed on first cycle: unknown cause 12541,5904860, 29 Dec 2017: pH failed at cycle 29 0692,5905073, 17 Jan 2017: pH sensor failed from outset, data are bad 12361,5905100, 18 Apr 2017: pH sensor is bad from outset. 12363,5905102, 30 May 2017: pH sensor went bad on Cy 5 12369,5905106, 26 Sept 2017: bad Aanderaa optode oxygen sensor on first profile, most likely a manufacturing defect 12729,5905133, 19 Sep 2017: pH failure upon deployment: As I've noted before, we see 3 different failure mechanisms based on diagnostic data that comes back. The fixes we've implemented might cure the first two, which are most common. The pH on float 12729 failed with the 3rd, more rare mechanism. Turns out there was some engineering data in the dock tests that indicated this might occur, but no one noticed. We're reviewing all the dock test data for other floats to verify they don't have this issue. When we were producing this batch of pH sensors, we had to scrap half a dozen or so with similar problems. It was due to a production short cut that the technicians put in place to speed things up. 12729,5905133, 26 Sep 2017: bad pH sensor on first profile, most likely an assembly defect 12745,5905132, 10 Oct 2017: pH sensor problems 12733,5905131, 8 Oct 2017 - box hit the water a bit hard on deployment 0690,5905071, 16 Jan 2018 - has bad pH sensor 0693,5905074, 12 Jan 2018: Navis float CTD failed, which means complete failure of float after the second profile. Failure occurred during parking phase. This float also had excess grease around the end cap prior to deployment and caused concern on part of I. Rosso, but deployment was OKed by D. Swift. 12757,5905366, 06 Feb 2018: pH and nitrate sensors are bad. float got knowcke around on the Agulhas 0688,5904846, 16 Jan 2018: pH sensor failed on first profile